OpenStudy (anonymous):

Please Help!. 1. Give an example of a whole number that is not a natural number. __________ 2. Give an example of an integer that is not a whole number. __________ 3. Give an example of a rational number that is not an integer. __________ 4. Give an example of an irrational number. ________ 5. Give an example of a complex number that is not an integer. __________ 6. Give an example of two integers whose sum is a natural number. ____________________ 7. Give an example of two rational numbers whose sum is an integer. ____________________

OpenStudy (anonymous):

number 1 could be 0

OpenStudy (anonymous):

yeah, and 2 cpuld be -1, right?

OpenStudy (anonymous):

Not sure what you mean. Try to Google your questions.

OpenStudy (anonymous):

I did, it didn't work

OpenStudy (anonymous):

Number 2 is any negative number

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

A negative number. Below zero. -3

OpenStudy (anonymous):

no, I know. I was saying do you know the answer to number 3

OpenStudy (anonymous):

Nember 3 is an improper fraction. The bottom number is not a facto of the top number. 4/3, 2/1, 6/5

OpenStudy (anonymous):

4 would be pi right?

OpenStudy (anonymous):

yes. 3.14159...

OpenStudy (anonymous):

5?

OpenStudy (anonymous):

I'm not sure about 5

OpenStudy (anonymous):

6?

OpenStudy (anonymous):

57 and 59

OpenStudy (anonymous):

I'm not sure about 7

OpenStudy (anonymous):

ok, thanks

OpenStudy (solomonzelman):

.5 + 1.5 I suppose, no?

OpenStudy (anonymous):

For which one?

OpenStudy (solomonzelman):

For number 7.

OpenStudy (anonymous):

okay, and 6?

OpenStudy (solomonzelman):

Any positive integer is a natural number from 1,2,3,4...... or even with negative interger \[-4+5,~~~~~~~~~or~~~~~~~~~3+1,~~~~~~or~~~~~~~~2+8\] anything almost.

OpenStudy (solomonzelman):

1000 of examples can be made up.

OpenStudy (anonymous):

Thanks, And 5?

OpenStudy (solomonzelman):

\[2i \sqrt{7}\] more?

OpenStudy (anonymous):

yeah, one sec.

OpenStudy (solomonzelman):

(I gave you an imaginary number) you know that \[i=\sqrt{-1}\]?

OpenStudy (anonymous):

8. Give an example of two complex numbers whose sum is a real number. ____________________ 9. Give an example of two rational numbers whose sum is a rational number. ____________________ 10. Give an example of two irrational numbers whose sum is an irrational number. ____________________ 11. Give an example of two integers whose product is a natural number. ____________________ 12. Give an example of two rational numbers whose product is an integer. ____________________ 13. Give an example of two complex numbers whose product is a real number. ____________________ 14. Give an example of two rational numbers whose product is a rational number. ____________________ 15. Give an example of two irrational numbers whose product is an irrational number. ____________________ 16. Give an example of two trinomials whose sum is a binomial. ________________________________________ 17. Give an example of two binomials whose sum is a trinomial. ________________________________________ 18. Give an example of two binomials whose sum is a binomial. ________________________________________ 19. Give an example of two binomials whose product is a binomial. ________________________________________ 20. Give an example of two polynomials whose product is a trinomial

OpenStudy (anonymous):

I am able to do the lessons by myself, but when it comes to these. the questions really confuse me.

OpenStudy (solomonzelman):

I am not assuming anything, I am going according to a definition of a "complex number" read this http://www.mathsisfun.com/definitions/complex-number.html you meant something else perhaps by saying complex number.

OpenStudy (anonymous):

Yeah, I get the whole imaginary number thing, but I just wanted to make sure I was answering the right way. Can you help me with the rest?

OpenStudy (solomonzelman):

But I can't think of a sum f 2 COMPLEX NUMBERS as you say, that would equal to a real number.

OpenStudy (solomonzelman):

Do you call\[2i+(-i)\] addition, then here you go!

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

I'll do number 9, and then I have to go.

OpenStudy (solomonzelman):

\[\frac{1}{3}+\frac{1}{2}\]

OpenStudy (anonymous):

okay, thank you!

OpenStudy (solomonzelman):

\[\frac{1}{3}+\frac{1}{2}=\frac{2}{6}+\frac{3}{6}=\frac{5}{6}\]

OpenStudy (solomonzelman):

repeating decimal, rational.

OpenStudy (solomonzelman):

Good luck with the rest, bye!